# Targeted Immunisation Thresholds for the Contact Process on Power-Law Trees

### v1

@@ 6.2 Small truncation threshold
For small $\rho$, the “small” truncation threshold $\mathbbm{t}=\rho\lambda^{-2}\log^{2}\frac{1}{\lambda}$ is no more sufficient to pass the infection from a big vertex to another big vertex at distance $c\log(1/\lambda)$ apart, and the previous survival strategy fails, but of course this is not sufficient to prove extinction of the epidemic.
@@
Observe that for large truncation threshold, a typical environment ensuring a high survival probability for the contact process was to consider the root being a big vertex, and other big vertices being found at distance roughly $c\log(1/\lambda)$ apart. In this section, we will in particular want to show that for a small truncation threshold, such an environment provides only a small survival probability. However, the probability of having the root and one direct neighbour [^4] Similarly one should still upper bound the survival probability if the root has 5 big neighbours, or any given constant number of big neighbours. of the root be big vertices is polynomially small in $\lambda$, so in order to obtain our main result we also should upper bound the survival provability of the infection on such environments.
Figure 1: We explore the ball of one (red) vertex of high degree, discovering another vertex of high degree and exploring its ball. Thus $|B|=2$ in this exploration.
@@
Inspired by these heuristics, we consider $\delta>0$ some parameter, with value to be determined later. We consider $S$ the smallest subtree containing the root such that any vertex $x\notin S$ with $\operatorname{d}(x,S)<\delta\log(1/\lambda)$ is small in the sense it has degree less than $1/8\lambda^{2}$. We also denote by $B$ the set containing the root and all other big vertices in $S$, and
$S^{+}=B(S,\lceil\delta\log(1/\lambda)\rceil)=B(B,\lceil\delta\log(1/\lambda)\rceil),$
@@
the $\lceil\delta\log(1/\lambda)\rceil$-neighbourhood of $S$, subtree containing $S$ and all vertices at distance less than $\lceil\delta\log(1/\lambda)\rceil$ from $S$. Note this tree $S^{+}$ contains no big vertex outside of $S$. Note also that we have
$|S|\leq|B|\lceil\delta\log(1/\lambda)\rceil,$
@@
as we can construct $S$ by starting from the root and sequentially adding each big vertex and a path of length at most $\lceil\delta\log(1/\lambda)\rceil$ linking it to the existing tree.
These trees are not necessarily finite, but we will call “good environment” an environment that satisfies
$|B|\leq\frac{\delta}{1000\rho},\qquad|S^{+}|\leq\frac{\delta}{1000\rho}\mathbbm{t}^{1+\delta\log\mu}.$
Note this definition depends on the value of $\delta$ and $\rho$. Lemma 6.3 below bounds the probability of observing a bad environment, while we will further bound the survival probability of the infection conditionally on any given good environment.
@@
First, by the monotonicity of the graphical construction it suffices to show the result on the tree with maximal root degree $\operatorname{d}(o)=\mathbbm{t}-1$. We then explore the ball around this root of radius $\left\lceil\delta\log\frac{1}{\lambda}\right\rceil$ for other big vertices – with degree at least $\frac{1}{8\lambda^{2}}$. Continue iteratively to explore the ball of radius $\left\lceil\delta\log\frac{1}{\lambda}\right\rceil$ around every big vertex until this construction terminates and we have explored the whole $S$, revealing the big vertices and paths between them of length at most $\left\lceil\delta\log\frac{1}{\lambda}\right\rceil$. To obtain the desired bounds on $|B|$ and $|S^{+}|$, we combine two ingredients:
  • We first control the size of the ball of radius $\left\lceil\delta\log\frac{1}{\lambda}\right\rceil$ around a given big vertex. We can upper bound the expected number of vertices in this ball by
    $\mathbbm{t}\sum_{j=1}^{\left\lceil\delta\log\frac{1}{\lambda}\right\rceil}\mu^{j-1}\leq\mathbbm{t}^{1+\frac{\log(1/\lambda)}{\log\mathbbm{t}}\delta\log\mu}/(1-1/\mu).$
    When $\lambda\to 0$, we have $\log(1/\lambda)/{\log\mathbbm{t}}\to 1/2<1$, so it is unlikely to observe more than $\mathbbm{t}^{1+\delta\log\mu}$ vertices in this ball. Actually, using [strayappendix, Theorem 1.1] with $n=\left\lceil\delta\log\frac{1}{\lambda}\right\rceil$, we obtain that the probability that the boundary of this ball contains more than $\mathbbm{t}^{1+\delta\log\mu}$ vertices decays like a stretched exponential in $\lambda$. By the union bound, the ball is then smaller than
    $\mathbbm{t}^{1+\delta\log\mu}\left\lceil\delta\log\frac{1}{\lambda}\right\rceil=\mathbbm{t}^{1+(\delta+o(1))\log\mu}$
    with a probability that’s also stretched exponential in $\lambda$, and so in particular a probability that is smaller than any polynomial in $\lambda$. By the union bound again, we get that with a probability which is still smaller than any polynomial in $\lambda$, the union of the balls of radius $\left\lceil\delta\log\frac{1}{\lambda}\right\rceil$ around the first $\frac{\delta}{1000\rho}$ explored big vertices contain less than $\frac{\delta}{1000\rho}\mathbbm{t}^{1+\delta\log\mu}$ vertices.
  • It remains to bound $|B|$ by $\frac{\delta}{1000\rho}$. To this end, it suffices to check that above $\frac{\delta}{1000\rho}\mathbbm{t}^{1+\delta\log\mu}$ explored vertices, it is likely to find less than $\frac{\delta}{1000\rho}$ big vertices. Note that for small enough $\lambda$ each explored vertex is big with probability
    $\mathbb{P}\left(D\geq\frac{1}{8\lambda^{2}}\right)<\mathbb{P}\left(X\geq\frac{1}{8\lambda^{2}}\right)<A\lambda^{2\tau-4}$
    and hence in $\left\lfloor\frac{\delta}{1000\rho}\right\rfloor\mathbbm{t}^{1+\delta\log\mu}$ explored vertices we find fewer than $\left\lfloor\frac{\delta}{1000\rho}\right\rfloor$ large vertices with probability, using the usual Chernoff bound, bounded by
    $\mathbb{P}\left(\operatorname{Bin}\left(\left\lfloor\frac{\delta}{1000\rho}\right\rfloor\mathbbm{t}^{1+\delta\log\mu},A\lambda^{2\tau-4}\right)\geq\left\lfloor\frac{\delta}{1000\rho}\right\rfloor\right)$
    $\leq\mathbb{P}\left(\operatorname{Pois}\left(\left\lfloor\frac{\delta}{1000\rho}\right\rfloor\rho^{1+\delta\log\mu}A\lambda^{2\tau-6-2\delta\log\mu}\log^{2+2\delta\log\mu}\frac{1}{\lambda}\right)\geq\left\lfloor\frac{\delta}{1000\rho}\right\rfloor\right)$
    $\leq\lambda^{\varphi\left\lfloor\frac{\delta}{1000\rho}\right\rfloor}\exp\left(\left\lfloor\frac{\delta}{1000\rho}\right\rfloor\rho^{1+\delta\log\mu}A\lambda^{2\tau-6-2\delta\log\mu-\varphi}\log^{2+2\delta\log\mu}\frac{1}{\lambda}\right)\sim\lambda^{\varphi\left\lfloor\frac{\delta}{1000\rho}\right\rfloor}\leq\lambda^{C}$
    where we have set parameter $\varphi={1000\rho C}/{\delta}$ so that by taking $\delta$ small and then $\rho$ small enough we guarantee $02\tau-6-2\delta\log\mu-\varphi>$ 0.
Taking $C>0$ arbitrary as in our main result, we thus can use Lemma 6.3 to fix $\delta$ and $\rho$ such that we have a good environment with failure probability bounded by $\lambda^{C}$. From now on, we work conditionally on a good environment, and aim to upper bound the survival probability of the infection in that case. We will more precisely bound the probability of ever hitting a vertex of $S^{+}$ at maximal distance $\lceil\delta\log(1/\lambda)\rceil$ from $S$.
We denote by $\mathscr{T}(S,B)\geq S$ the infinite tree containing $S$ in which
$\forall v\in B\quad$ $\operatorname{d}(v)=\mathbbm{t},$
$\forall v\in V(\mathscr{T})\setminus B\quad$ $\operatorname{d}(v)=\left\lfloor\frac{1}{8\lambda^{2}}\right\rfloor.$
Write also $\partial S$ for the set of edges of $\mathscr{T}(S,B)$ at the boundary of $S$, which has
$|\partial S|=|S|\left\lfloor\frac{1}{8\lambda^{2}}\right\rfloor-|S|+1+|B|\left(\mathbbm{t}-\left\lfloor\frac{1}{8\lambda^{2}}\right\rfloor\right)\geq\frac{|S|}{10\lambda^{2}},$
as well as
$|\partial S|\leq|B|\mathbbm{t}+(|S|-|B|)\frac{1}{8\lambda^{2}}\leq|B|(\mathbbm{t}+o(\mathbbm{t}))\leq\frac{\delta\mathbbm{t}}{900\rho}=\frac{\delta\log^{2}\frac{1}{\lambda}}{900\lambda^{2}}.$ (13)
Lemma 6.4.
Consider the neighbourhood infection process for the tree, seen in Lemma 5.1
$H_{t}:=\sum_{v\notin S}\xi_{t}(v)(2\lambda)^{\operatorname{d}(v,S)-1}$
but now on this new tree $\mathscr{T}(S,B)$ with $B$ nonempty. From any initial configuration $\xi_{0}$ with $H_{0}<15\lambda|\partial S|$, and for $\lambda$ sufficiently small, we find
@@
$\mathbb{P}_{\xi_{0}}\left(\sup_{t\leq 7^{\lambda|\partial S|}}H_{t}>20\lambda|\partial S|\right)\leq\exp\left(-\frac{1}{100\lambda}\right).$

### v2

@@ 6.2 Small truncation threshold
For small $\rho$, the “small” truncation threshold $\mathbbm{t}=\rho\lambda^{-2}\log^{2}(\nicefrac{{1}}{{\lambda}})$ is no more sufficient to pass the infection from a big vertex to another big vertex at distance $c\log(\nicefrac{{1}}{{\lambda}})$ apart, and the previous survival strategy fails, but of course this is not sufficient to prove extinction of the epidemic.
@@
Observe that for large truncation threshold, a typical environment ensuring a high survival probability for the contact process was to consider the root being a big vertex, and other big vertices being found at distance roughly $c\log(\nicefrac{{1}}{{\lambda}})$ apart. In this section, we will in particular want to show that for a small truncation threshold, such an environment provides only a small survival probability. However, the probability of having the root and one direct neighbour of the root both big vertices is polynomially small in $\lambda$, so in order to obtain our main result we also should upper bound the survival provability of the infection on such an environment [^4] Similarly, the event that the root has 5 big neighbours – or any given constant number of big neighbours – is also polynomially unlikely, and so we also have to bound the survival probability on this event..
Figure 1: We explore the $\lceil\delta\log(\nicefrac{{1}}{{\lambda}})\rceil$-ball of one (big) vertex of high degree, discovering another vertex of high degree. Thus $|\mathcal{B}|=2$ in this exploration but the four highlighted vertices are the set $\mathcal{C}$, around each of which we explored a ball of radius $\lceil\delta\log(\nicefrac{{1}}{{\lambda}})\rceil$ and found no further big vertices. Note also $\mathcal{C}^{+}$ is the set of every vertex in this picture. In our proof, the whole set $\mathcal{C}$ acts as a big vertex for which we want to control the number of infections sent far away as in Subsection 5.2. Colour available online.
@@
Inspired by these heuristics, we consider $\delta>0$ some parameter, with value to be determined later. We consider the core $\mathcal{C}$ the smallest subtree containing the root such that any vertex $x\notin\mathcal{C}$ with $\operatorname{d}(x,\mathcal{C})<\delta\log(\nicefrac{{1}}{{\lambda}})$ is small in the sense it has degree less than $1/8\lambda^{2}$. We also denote by $\mathcal{B}$ the set containing the root and all other big vertices in $\mathcal{C}$, and
$\mathcal{C}^{+}=B(\mathcal{C},\lceil\delta\log(\nicefrac{{1}}{{\lambda}})\rceil),$
@@
the $\lceil\delta\log(\nicefrac{{1}}{{\lambda}})\rceil$-neighbourhood of $\mathcal{C}$. Note this tree $\mathcal{C}^{+}$ contains no big vertex outside of $\mathcal{C}$. Note also that we have
$|\mathcal{C}|\leq|\mathcal{B}|\lceil\delta\log(\nicefrac{{1}}{{\lambda}})\rceil,$
@@
as we can construct $\mathcal{C}$ by starting from the root and sequentially adding each big vertex and a path of length at most $\lceil\delta\log(\nicefrac{{1}}{{\lambda}})\rceil$ linking it to the existing tree.
Note that in this construction, nothing guarantees that $|\mathcal{B}|$, $|\mathcal{C}|$ or $|\mathcal{C}^{+}|$ are small or even finite, however part of the proof will be to show that with an appropriate choise of $\delta$ and $\rho$ and with probability $1-o(1)$ in the limit of small $\lambda$, these sizes can be controlled, and more precisely
$|\mathcal{B}|\leq\frac{\delta}{400\rho},\qquad|\mathcal{C}^{+}|\leq\frac{\delta}{400\rho}\mathbbm{t}^{1+\delta\log\mu}.$
An environment satisfying these restrictions will be called a “good environment”. Thus in a good environment we have fewer than a (large) constant number of large vertices, and then their average ball size is less than $\omega(1)$ factor larger than expectation. Lemma 6.3 below bounds the probability of observing a bad environment.
@@
We postpone the proof of this lemma to the end of the section, and continue with our main proof. Taking $C>0$ arbitrary as in our main result, we can use Lemma 6.3 to fix $\delta$ and $\rho$ such that we have a good environment with failure probability bounded by $\lambda^{C}$. From now on, we work conditionally on a good environment, and aim to upper bound the survival probability of the infection in that case. We will more precisely bound the probability of ever hitting a vertex of $\mathcal{C}^{+}$ at maximal distance $\lceil\delta\log(1/\lambda)\rceil$ from $\mathcal{C}$.
Heuristically, we follow the same guidelines as in Subsection 5.2, with the complication that the central big vertex of that section is replaced by the whole set $\mathcal{C}$. We denote by $\mathscr{T}(\mathcal{C},\mathcal{B})\geq\mathcal{C}$ the infinite tree containing $\mathcal{C}$ in which
$\forall v\in\mathcal{B}$ $\operatorname{d}(v)=\mathbbm{t},$
$\forall v\in V(\mathscr{T})\setminus\mathcal{B}$ $\operatorname{d}(v)=\left\lfloor\frac{1}{8\lambda^{2}}\right\rfloor.$
Write also $\partial\mathcal{C}$ for the set of edges of $\mathscr{T}(\mathcal{C},\mathcal{B})$ at the boundary of $\mathcal{C}$, which has
$|\partial\mathcal{C}|=|\mathcal{C}|\left\lfloor\frac{1}{8\lambda^{2}}\right\rfloor-|\mathcal{C}|+1+|\mathcal{B}|\left(\mathbbm{t}-\left\lfloor\frac{1}{8\lambda^{2}}\right\rfloor\right)\geq\frac{|\mathcal{C}|}{10\lambda^{2}},$
as well as
$|\partial\mathcal{C}|\leq|\mathcal{B}|\mathbbm{t}+(|\mathcal{C}|-|\mathcal{B}|)\frac{1}{8\lambda^{2}}\leq|\mathcal{B}|(\mathbbm{t}+o(\mathbbm{t}))\leq\frac{\delta\mathbbm{t}}{300\rho}=\frac{\delta\log^{2}\frac{1}{\lambda}}{300\lambda^{2}}.$ (12)
Similarly as in the proof of Lemma 5.1, we work on the tree $\mathscr{T}(\mathcal{C},\mathcal{B})$ and consider the exponentially weighted sum
$H_{t}:=\sum_{v\notin\mathcal{C}}\xi_{t}(v)(2\lambda)^{\operatorname{d}(v,\mathcal{C})-1},$
which controls the infection outside of $\mathcal{C}$, and has expected infinitesimal increase bounded by
[…]
So the process $H_{t}$ is likely to quickly go to 0 when $\mathcal{C}$ is healthy, and is likely to quickly go below $4\lambda|\partial\mathcal{C}|$ (or close to this value) when $\mathcal{C}$ is infected. The set $\mathcal{C}$ is also infected at rate bounded by $\lambda H_{t}$, however we cannot simply argue that it recovers at rate 1, and so here is a limit of the similarity with Lemma 5.1. Instead, we also have an infection process on $\mathcal{C}$
$U_{t}:=\sum_{v\in\mathcal{C}}\xi_{t}(v),$
with expected infinitesimal increase bounded by
$\frac{\operatorname{\mathbb{E}}\left[U_{t+\rm dt}-U_{t}|\mathcal{F}_{t}\right]}{{\rm d}t}\leq\lambda H_{t}-U_{t}+\lambda|\mathcal{C}|U_{t}$ (14)
using of course that the number of edges of the tree $\mathcal{C}$ is $|\mathcal{C}|-1\leq|\mathcal{C}|$. In our proof we have to study the dynamics of both processes $H$ and $U$, with the difficuly that the dynamics of any of these two processes depends on the other. In order to simplify the picture, we would like to get rid of the $H$-dependence of the dynamics of $U$. It is possible if we have an upper bound on $H$ which can serve as an upper bound for the dynamics of $U$, and to this end we will use the following lemma, which states that $H$ is likely to stay a long time below level $20\lambda|\partial\mathcal{C}|$.
Lemma 6.4.
For the process $H$ defined on the tree $\mathscr{T}(\mathcal{C},\mathcal{B})$ with $\mathcal{B}$ nonempty, from any initial configuration $\xi_{0}$ with $H_{0}<15\lambda|\partial\mathcal{C}|$, and for $\lambda$ sufficiently small, we find
$\mathbb{P}_{\xi_{0}}\left(\sup_{t\leq 7^{\lambda|\partial\mathcal{C}|}}H_{t}>20\lambda|\partial S|\right)\leq\exp\left(-\frac{1}{100\lambda}\right).$
We postpone also the proof of this lemma to the end of the section, and continue with an adaptation of Lemma 5.1 and Corollary 5.2 to our new settings.
Lemma 6.5.
Assuming $\lambda$ is sufficiently small, on the tree $\mathscr{T}(\mathcal{C},\mathcal{B})$ we have
$\operatorname{\mathbb{E}}_{\mathbbm{1}_{\mathcal{C}}}\left[\int_{0}^{\infty}\mathbbm{1}_{\xi_{t}(\mathcal{C})>0}\,{\rm d}t\right]\leq 3e^{200\lambda^{2}|\partial\mathcal{C}|}.$
@@
We introduce the sequences of stopping times $(\tau_{k})_{k\geq 1}$, $(R_{k})_{k\geq 1}$, $(S_{k})_{k\geq 1}$ and $(T_{k})_{k\geq 0}$ defined recursively by $T_{0}=0$ and for $k\geq 1:$
$\tau_{k}$ $:=\inf\{t\geq T_{k-1},H_{t}\geq 20\lambda|\partial\mathcal{C}|\},$
$R_{k}$ $:=\inf\{t\geq T_{k-1},\xi_{t}(\mathcal{C})=0\}\wedge\tau_{k},$
$S_{k}$ :=$:=\begin{cases}R_{k}&\text{if }R_{k}=\tau_{k}\\ \inf\{t\geq R_{k},\xi_{t}(\mathcal{C})>0\}&\text{if }R_{k}<\tau_{k},\end{cases}$
$T_{k}$ $:=\inf\{t\geq S_{k},H_{t}\leq 15\lambda|\partial\mathcal{C}|\}.$
In words, starting from a time $t=T_{k-1}$ with $H_{t}\leq 15\lambda|\partial\mathcal{C}|$, we wait until either $H$ goes above $20\lambda|\partial\mathcal{C}|$ or $\mathcal{C}$ recovers. If we first observe $H$ going above the threshold, we just wait until it goes down below $15\lambda|\partial\mathcal{C}|$ again (at time $T_{k}$). If on the contrary we first have recovery of $\mathcal{C}$, we first check whether it is a permanent recovery (in which case we have $S_{k}=T_{k}=+\infty$) and if not, after the reinfection of $\mathcal{C}$ we still wait until $H$ goes down below $15\lambda|\partial\mathcal{C}|$ at time $T_{k}$ to repeat the procedure. We claim that:
  1. (i)
    Conditionally on $T_{k-1}$ being finite, we have the upper bound
@@
    $\operatorname{\mathbb{E}}\left[(R_{k}-T_{k-1})+(T_{k}-S_{k})\Big{|}T_{k-1}<\infty\right]\leq 15+e^{40\lambda^{2}|\partial\mathcal{C}|}$
    for the average time $\mathcal{C}$ is infected on $[T_{k-1},T_{k}]$.
  2. (ii)
    Conditionally on $T_{k-1}$ being finite, we have a probability at least $\frac{2}{5}e^{-160\lambda^{2}|\partial\mathcal{C}|}$ of having $T_{k}$ infinite, actually by observing a permanent recovery of $\mathcal{C}$, leading to $S_{k}$ being infinite.
Combining the two claims provides the upper bound
$\operatorname{\mathbb{E}}\left[\int_{0}^{\infty}\mathbbm{1}_{\xi_{t}(\mathcal{C})>0}{\rm d}t\Bigg{|}\xi_{0}=\mathbbm{1}_{\mathcal{C}}\right]\leq\left(15+e^{40\lambda^{2}|\partial\mathcal{C}|}\right)\left(\frac{5}{2}e^{160\lambda^{2}|\partial\mathcal{C}|}\right),$
and thus the result for small $\lambda$.
To check the claim in $(i)$, we repeat the supermartingale argument in Lemma 5.1, but now with the process $11\log(1+H_{t}/15\lambda|\partial\mathcal{C}|)+\nicefrac{{t}}{{11}}$, which has nonnegative infinitesimal increase when $H$ is above $15\lambda|\partial\mathcal{C}|$, to deduce
$\operatorname{\mathbb{E}}\left[T_{k}-S_{k}\Big{|}T_{k-1}<\infty\right]\leq 11\log\left(1+\frac{\operatorname{\mathbb{E}}[H_{S_{k}}|T_{k-1}<\infty]}{15\lambda|\partial\mathcal{C}|}\right).$
Still working implicitly conditionally on $\{T_{k-1}<+\infty\}$, we bound the inner expectation by
$\begin{array}[]{rlclcl}\operatorname{\mathbb{E}}[H_{S_{k}}]&\leq&\operatorname{\mathbb{E}}[H_{S_{k}}\mathbbm{1}_{R_{k}=\tau_{k}}]&+&\operatorname{\mathbb{E}}[H_{S_{k}}\mathbbm{1}_{R_{k}<\tau_{k}}]&\\ &\leq&(1+20\lambda|\partial\mathcal{C}|)&+&20\lambda|\partial\mathcal{C}|&\leq 41\lambda|\partial\mathcal{C}|,\end{array}$
where in the second line we used that the jumps of $H$ are bounded by one to upper bound the process by $1+20\lambda|\partial\mathcal{C}|$ at time $\tau_{k}$, and in the second term the supermartingale property of the process $H$ when $\mathcal{C}$ is healthy which gives a nonpositive expected increment on the time interval $[R_{k},S_{k}]$. Thus
$\operatorname{\mathbb{E}}\left[T_{k}-S_{k}\Big{|}T_{k-1}<\infty\right]\leq 11\log(1+\nicefrac{{41}}{{15}})\leq 15.$
In order to bound $\operatorname{\mathbb{E}}[R_{k}-T_{k-1}|T_{k-1}<+\infty]$, it suffices to bound the recovery time of $\mathcal{C}$, or equivalently the hitting time of 0 for the process $U$, under the hypothesis that $H$ stays bounded by $20\lambda|\partial\mathcal{C}|$. Recalling (14), we can then bound its infinitesimal increase by $220\lambda^{2}|\partial\mathcal{C}|-U_{t}/$ 2, using that for small $\lambda$ we have $2\lambda|\mathcal{C}|\leq\frac{2\delta}{400\rho}\lambda\lceil\delta\log(1/\lambda)\rceil\leq 1$. Further though, this upper bound is just a birth-death chain on $\{0\}\cup[|\mathcal{C}|]$ which by detailed balance has a stationary distribution $\pi$ with
$\forall k\in[|\mathcal{C}|]\quad 40\lambda^{2}|\partial\mathcal{C}|\pi(k-1)=k\pi(k)\implies\frac{\pi(k)}{\pi(0)}=\frac{\left(40\lambda^{2}|\partial\mathcal{C}|\right)^{k}}{k!}$
$\pi(0)\frac{\left(40\lambda^{2}|\partial\mathcal{C}|\right)^{k}}{2(k-1)!}$
$\frac{1}{\pi(0)}\sum_{k=1}^{|\mathcal{C}|}\frac{2(k-1)!}{\left(40\lambda^{2}|\partial\mathcal{C}|\right)^{k}}$ $=\sum_{k=1}^{|\mathcal{C}|}\frac{2(k-1)!}{\left(40\lambda^{2}|\partial\mathcal{C}|\right)^{k}}\sum_{j=1}^{|\mathcal{C}|}\frac{\left(40\lambda^{2}|\partial\mathcal{C}|\right)^{j}}{j!}$
$\leq|\mathcal{C}|\frac{2(|\mathcal{C}|-1)!}{\left(40\lambda^{2}|\partial\mathcal{C}|\right)^{|\mathcal{C}|}}\sum_{j=1}^{\infty}\frac{\left(40\lambda^{2}|\partial\mathcal{C}|\right)^{j}}{j!}\leq\frac{2(|\mathcal{C}|)!e^{40\lambda^{2}|\partial\mathcal{C}|}}{\left(40\lambda^{2}|\partial\mathcal{C}|\right)^{|\mathcal{C}|}}.$
Then, use that $|\partial\mathcal{C}|\geq\frac{|\mathcal{C}|}{10\lambda^{2}}$ and $n!\leq\frac{n^{n+1}}{e^{n-1}}$ to bound this hitting time by
$\frac{2(|\mathcal{C}|)!e^{40\lambda^{2}|\partial\mathcal{C}|}}{\left(40\lambda^{2}|\partial\mathcal{C}|\right)^{|\mathcal{C}|}}\leq\frac{2(|\mathcal{C}|)!e^{40\lambda^{2}|\partial\mathcal{C}|}}{\left(4|\mathcal{C}|\right)^{|\mathcal{C}|}}\leq 2|\mathcal{C}|e{(4e)}^{-|\mathcal{C}|}e^{40\lambda^{2}|\partial\mathcal{C}|}\leq e^{40\lambda^{2}|\partial\mathcal{C}|},$
and we have shown the claim in $(i)$.
For the second point, we first note that we have $R_{k}<\tau_{k}$ with probability at least $4/5$. Indeed, by Lemma 6.4 and Markov’s inequality, we will have hitting of $U=0$ before time $7^{\lambda|\partial\mathcal{C}|}$, and hitting of $H>20\lambda|\partial\mathcal{C}|$ after time $7^{\lambda|\partial\mathcal{C}|}$, with probability at least
$1-\exp\left(-\frac{1}{100\lambda}\right)-\exp\left(\left(40\lambda-\log 7\right)\lambda|\partial\mathcal{C}|\right)\geq 1-\exp\left(-\frac{1}{5\lambda}\right)\geq\frac{4}{5}.$
Now, conditionally on $R_{k}$ being finite and smaller than $\tau_{k}$, we can condition on $\xi_{R_{k}}$ (a configuration which has in particular $H_{R_{k}}\leq 20\lambda|\partial\mathcal{C}|$), and observe that $(H_{t\wedge S_{k}}e^{(t\wedge S_{k})/4})_{t\geq R_{k}}$ is a supermartingale to deduce
$\mathbb{P}\left(\int_{R_{k}}^{S_{k}}H_{t}>160\lambda|\partial\mathcal{C}|\,\Big{|}\,\xi_{R_{k}}\right)\leq\frac{\operatorname{\mathbb{E}}\left[\int_{R_{k}}^{S_{k}}H_{t}{\rm d}t\,\Big{|}\,\xi_{R_{k}}\right]}{160\lambda|\partial\mathcal{C}|}\leq\frac{20\lambda|\partial\mathcal{C}|\int_{0}^{\infty}e^{-t/4}{\rm d}t}{160\lambda|\partial\mathcal{C}|}=\frac{1}{2}.$
If this limit is not exceeded, the total rate of reinfection of $\mathcal{C}$ is then bounded by $\lambda\cdot 160\lambda|\partial\mathcal{C}|$, and so we have found that the recovery of $\mathcal{C}$ is permanent with probability at least $\frac{1}{2}e^{-160\lambda^{2}|\partial\mathcal{C}|}$, which allows to conclude the claim in $(ii)$, and thus our proof. ∎
Corollary 6.6.
On the tree $\mathscr{T}(\mathcal{C},\mathcal{B})$, generalise our CP-BRW notation (4) so that $m(\mathcal{C},v)$ denotes the mean absorption at $v$ with the good environment $\mathcal{C}$ initially fully infected. When the target sink $v$ is at distance $\lceil\delta\log(1/\lambda)\rceil$ from $\mathcal{C}$, we find
$m(\mathcal{C},v)\leq e^{-\frac{1}{4}\delta\log^{2}\frac{1}{\lambda}}.$